A constant Force, F of magnitude 20N pulled a light cord wrapped around a pulley to lift a bucket of mass 1.53kg. The radius and the moment of inertia of the pulley are 0.33m and 0.385kgm^2 respectively. If the frictional torque is 1.10Nm calculate the angular acceleration of the pulley.
1
Expert's answer
2018-07-26T11:27:09-0400
The angular acceleration is
α=T/I.
T=(F-mg)r-T_fr
So,
α=((F-mg)r-T_fr)/I=((20-1.53∙9.8)0.33-1.1)/0.385=1.43 rad/s^2 .